# Quantum criticality and universality in stationary state of long-range Kitaev model

### v1

@@ Abstract
We investigate the signature of quantum criticality in the long-time stationary state of the long-range Kitaev chain by performing various quench protocols. In this model, the pairing interaction decays with distance according to a power law with exponent $\alpha$. Using quantum information-theoretic measures, such as mutual information and logarithmic negativity, we show that, irrespective of the values of $\alpha$, critical-to-critical quench displays quantum criticality even in the stationary state. Remarkably, in the presence of long-range pairing interactions, where fermionic correlators decay algebraically even at non-critical points, signature of quantum criticality persists in the stationary state. Furthermore, the effective central charge, calculated from both mutual information and logarithmic negativity of stationary state following a critical-to-critical quench, agrees with the central charge of the corresponding ground states for both $\alpha=0$ and $\alpha=2$. Therefore, information of the universality class can be inferred from the stationary state.
@@ I Introduction
Understanding quantum phase transitions (QPTs) has been a central topic of interest over the past few decades due to its significant implications for understanding the collective behavior of many-body systems [1]. QPTs have been both theoretically and experimentally studied extensively in various systems, including quantum spin chains [2; 3; 4; 5; 6; 7; 1; 8; 9; 10; 11; 12], Bose-Einstein condensates [13; 14; 15; 16; 17; 18; 19; 20], and strongly correlated electron systems [21; 22; 23; 24; 25]. QPT occurs at specific values of the parameter(s) present in Hamiltonian, called critical points where the system exhibits scale invariance. A universal behavior ensues due to the scale invariance, which can be classified into certain universality classes based on the space dimension and symmetry of the order parameter[26; 27; 28; 29]. This universality is studied using conformal field theories, and the corresponding central charge value is commonly used to identify the underlying universality class [30; 31; 32]. For instance, systems belonging to the Ising universality class are described by a minimal model with a central charge $c=1/2$, while the Luttinger liquid universality class corresponds to a conformal bosonic theory with $c=1$ [33; 34]. Note that at quantum critical points, the development of long-range correlations leads to an algebraic decay of correlators over distance, in contrast to the exponential decay typically observed at noncritical points in Hamiltonians with only short-range interactions [35; 36; 37; 38]. This is often used to identify critical points.
@@
While the existence of quantum criticality and long-range correlations in the ground states of many-body quantum systems are widely studied, their presence in stationary states are not yet well explored. In Ref. [39], it is shown that when two sides of a 1D non-interacting fermionic chain with nearest-neighbor hopping are prepared at different temperatures, the mutual information scales logarithmically with subsystem size in the steady state, indicating the presence of long-range correlations. Another example of the survival of quantum criticality in the long-time stationary state is shown in the anisotropic XY chain under a sudden quench protocol [40]. Using quantum information measures such as mutual information and logarithmic(log-) negativity, it is shown that when both the pre- and post-quench parameters of the Hamiltonian are set at the critical point, both mutual information and log-negativity exhibit a peak, indicating the presence of criticality in the stationary state. In contrast, the peak in mutual information and log-negativity vanishes for other quench protocols when either/both pre- and post-quench parameters differ from critical values. The signature of quantum criticality in the stationary state is attributed to a change in correlation pattern from exponential to algebraic decay, which was captured by mutual information and log-negativity.
@@
A natural question follows about the scenario when correlators decay algebraically, even at noncritical points? Long-range interacting systems provide an ideal platform to explore this question as fermionic correlators have been shown to decay algebraically over distance, even when the Hamiltonian is gapped [41; 42; 43; 44], i.e. at noncritical points. These long-range interacting systems have been studied in different contexts both theoretically and experimentally in atomic, molecular, and optical lattice systems [45; 46; 47; 48; 49; 50; 51; 52; 53; 54; 55; 56; 57; 58; 59; 60; 61]. We take the 1D long-range Kitaev model with a pairing term that decays with distance as a power law $\propto l^{-\alpha}$ with $\alpha$ being the exponent[44]. From an experimental perspective, this model is particularly relevant as it is closely related to the Ising model with tunable long-range interactions, which can be experimentally realized using trapped ion setups [53; 54; 55; 62; 58; 57]. In the ground state, this model undergoes an exotic transition from the Ising type universality class, observed for $\alpha>3/2$, to a Luttinger liquid universality class at $\alpha=0$.
@@
In this paper, we investigate whether it is possible to capture quantum criticality in the stationary state when fermionic correlators decay algebraically even at non-critical points. Additionally, we examine whether the stationary state can be described by the same universality class as the ground state by analyzing the central charge obtained via scaling of mutual information and log-negativity for various quench protocols. Specifically, we consider three values of $\alpha$ corresponding to different universality classes observed in the ground state: first, corresponding to the Ising universality class ($\alpha=2$); second, the Luttinger liquid universality class ($\alpha=0$); and a third value ($\alpha=1$) where there is no universality [44].
@@ II The model and ground state phase transition
where $f_{j}(f_{j}^{\dagger})$ is the fermionic annihilation (creation) operator at site $j$, satisfying the canonical anticommutation relations $\{f_{i},f_{j}^{\dagger}\}=\delta_{ij}$, $\{f_{i},f_{j}\}=\{f_{i}^{\dagger},f_{j}^{\dagger}\}=0$. The parameters $t$ and $\mu$ represent the tunneling rate between two neighboring sites and the chemical potential, respectively. We set $12t=$ 1 in the rest of the paper. The parameter, $\Delta$ denotes the strength of the fermion $p$-wave pairing interaction, while its range is governed by the exponent $\alpha\in[0,\infty)$. The two limits, i.e. $\alpha=0$ and $\alpha\to\infty$ correspond to all-to-all interaction with equal strength and nearest-neighbour interaction, respectively. We consider the anti-periodic boundary condition throughout the paper. The Hamiltonian in Eq. 1 is exactly solvable. To see this, we first perform a Fourier transformation to transform Eq. 1 in the momentum space as
@@
where $g_{\alpha}(k)=\sum_{l=1}^{N-1}\frac{\sin(kl)}{l^{\alpha}}$ and $k_{n}=\frac{2\pi}{N}\left(n+1/2\right)$. In the large $N$ limit, $g_{\alpha}(k)$ takes the form
@@
The function $\rm Li_{\alpha}(z)$ represents the polylogarithm of the complex variable $z$ of order $\alpha$. The Hamiltonian in Eq. 2 can further be cast to a diagonal form by performing the following Bogoliubov transformation
@@
with $\theta$ being the Bogoliubov angle and is defined as
@@
Equation 6 signifies that each modes are independent implying the integrability of the system. In the limit $\alpha\to\infty$, only nearest neighbour terms contribute to the sum in $g_{\alpha}(k)$, resulting in $g_{\infty}(k)=\sin(k)$. The Hamiltonian in Eq. 1 in this limit coincides with the Hamiltonian of the $XY$ model obtained via the Jordan-Wigner transformation. For the $XY$ model, a ground state phase transition occurs from the gapped ferromagnetic ordered phase to the gapped paramagnetic disordered phase with increasing the chemical potential $|\mu|$ with the two critical gapless points at $\mu=\pm 1$ [44]. The spectrum in Eq. 7 continues to be gapless at $\mu=\pm 1$ even with the smaller values of $\alpha$, i.e., in the presence of long-range interactions as long as $\alpha>1$. However, the phase diagram changes for $\alpha<1$. This regime is referred to as “strong” long-range regime since $\alpha<d$, with $d$ representing the dimension of the system. In comparison to $\alpha>1$, one significant change is that $\mu=-1$ is no longer a critical point, and the phase diagram is no longer symmetric across the line $\mu=0$ [44; 42].
Table 1: Summary of the effective central charge $c_{\rm eff}$ for different values of $\alpha$ and $\mu$ for ground state of long-range Kitaev Model [44; 42; 63; 64].
@@
A good probe to identify these critical points is the von Neumann entropy, $S_{vN}$, expressed as $S_{vN}=-{\text{Tr}}\left(\rho_{L}\ln\rho_{L}\right)$, where $\rho_{L}$ is the reduced density matrix of the subsystem of length $L$. At these critical points, where the spectral gap closes, $S_{vN}$ for short-range interacting system scales as [65; 66; 67; 68],
@@
where $c$ is the central charge of the underlying conformal field theory (CFT), $N$ is the total system size and $c^{\prime}$ is a non-universal constant. In fact, Eq. 8 is also valid for all values of $\alpha$ and $\mu$, but the central charge $c$ is now replaced with an effective central charge $c_{\rm eff}$ [44; 63]. The values of the $c_{\rm eff}$ for ground state are summarized in Table-1. It is worth mentioning that $\alpha=0$ signifies all-to-all coupling between the sites and falls under the Tomonaga-Luttinger liquid universality class [69].
@@ III Methods
In last section, we summarized criticality seen in ground state of LRK model and (effective) central charge of the underlying CFT. A natural question then arises whether this criticality survives when system undergoes out of equilibrium dynamics? In other words, in a stationary state can we find any signature of quantum criticality? A common way to achieve the non-equilibrium dynamics, is through sudden quench where the local or global parameters of the system are suddenly changed [70]. As a consequence, the ground state of the pre-quench Hamiltonian is no longer a ground state of the post-quench Hamiltonian but a superposition of it’s eigenstates. The long-time stationary state of the LRK model under sudden quench is the system of interest. To probe the criticality, much like ground state, entanglement entropy would be our first choice. However, as the long-time stationary state typically involves highly excited states, $S_{vN}$ exhibits volume-law scaling eclipsing the logarithmic dependence on subsystem size which may be present due to criticality. So, to extract these logarithmic correlations, we consider quantum information theoretic measures such as mutual information and log-negativity. For ground state, mutual information has already been established to capture the criticality in fermionic systems [71].
Mutual information: The mutual information between two subsystems $A_{1}$ and $A_{2}$ is defined as
@@
where $S_{vN}^{A_{1}}(S_{vN}^{A_{2}})$ is the von Neumann entropy of the subsystem $A_{1}(A_{2})$ and $S_{vN}^{A_{1}\cup A_{2}}$ is the von Neumann entropy of $A_{1}\cup A_{2}$. Throughout the paper, we consider the subsystem sizes $|A_{1}|=|A_{2}|=L$. The mutual information between two subsystems measures the total amount of information (both classical and quantum) that one subsystem contains about the other [72]. Furthermore, if $S_{vN}$ of the long time stationary state contains both the volume and logarithmic term, the mutual information should scale as $I_{A_{1}:A_{2}}\sim b\ln L$. Thus, if the long-range correlations are present in the stationary state, then the mutual information should behave as [73; 74; 75; 76; 39; 40]
@@
Log-negativity: Similar to mutual information, log-negativity, an entanglement monotone, also quantifies the non-local correlation between two subsystems. But unlike mutual information, it captures only quantum correlations [77; 78; 79; 80]. It is defined as [78; 77]
@@
where $c$ is the central charge. For the $XY$ model, during a critical-to-critical quench, the log negativity of the long-time stationary state scales as $\xi_{A_{1}:A_{2}}=\frac{1}{8}\ln L$, where $L_{1}=L_{2}=L$ [40]. The logarithmic divergence of $\xi_{A_{1}:A_{2}}$ is consistent with the critical ground state, as the $XY$ model belongs to the Ising universality class with $c=1/2$. Based on this, we assume that if there is any signature of long-range quantum correlations in the long time stationary state, the log negativity would exhibit the following scaling behaviour:
@@
where $c_{\rm eff}^{N}$ represents the effective central charge of the stationary state extracted using the log negativity.
Time-evolution: For non-interacting free-fermionic chains, owing to the simple relationship between the eigenvalues of reduced density matrix and two-point correlation function, we can calculate the entanglement entropy more efficiently by defining correlation matrix $W$ as,
@@
$S=-\frac{1}{2}{\text{Tr}}\left[(I-W)\log(I-W)+W\log W\right].$ (15)
At any finite time $t$, the correlation functions are given by $C_{nm}(t)=\langle f_{n}^{\dagger}(t)f_{m}(t)\rangle$ and $F_{nm}(t)=\langle f_{n}(t)f_{m}(t)\rangle$, where the expectation values are taken with respect to the time evolved state obtained under the operation of the post-quench Hamiltonian. The two time-dependent correlation functions, $C_{nm}(t)$ and $F_{nm}(t)$ are expressed as [83]
$\langle f_{j}(t)f_{j+l}(t)\rangle$ $=\frac{1}{2N}\sum_{n=0}^{N-1}e^{ik_{n}l}\bigg{[}-\sin\left(2\lambda^{f}_{\alpha}(k_{n})t\right)\sin\left(2\delta\theta_{k_{n}}\right)$ (16)
$\quad+i\bigg{(}\sin\left(2\theta_{k_{n}}\right)\cos\left(2\delta\theta_{k_{n}}\right)$
$\qquad-\cos\left(2\lambda^{f}_{\alpha}(k_{n})t\right)\sin\left(2\delta\theta_{k_{n}}\right)\cos\left(2\theta_{k_{n}}\right)\bigg{)}\bigg{]}$
$\langle f_{j}^{\dagger}(t)f_{j+l}(t)\rangle$ $=\frac{1}{2N}\sum_{n=0}^{N-1}e^{-ik_{n}l}\bigg{[}1-\cos\left(2\theta_{k_{n}}\right)\cos\left(2\delta\theta_{k_{n}}\right)$
$\quad-\sin\left(2\theta_{k_{n}}\right)\sin\left(2\delta\theta_{k_{n}}\right)\cos\left(2\lambda^{f}_{\alpha}(k_{n})t\right)\bigg{]},$
where $\delta\theta_{k_{n}}=\theta^{f}_{k_{n}}-\theta^{i}_{k_{n}}$ is the difference between pre- and post-quench Bogoliubov angles defined in Eq. 5 and $\lambda^{f}_{\alpha}(k_{n})$ is defined by Eq. 7 for post-quench Hamiltonian.
In the limit $t\to\infty$, the time dependent sine and cosine functions become highly oscillatory and the respective sum would tend to zero. Using this approximation, the expression for the correlation functions in the long-time stationary state is given by,
@@
where $\langle f_{j}^{\dagger}f_{j+l}\rangle_{\text{st}}$ and $\langle f_{j}f_{j+l}\rangle_{\text{st}}$ are two point correlation and two-point anomalous correlation function for stationary state. We utilize the above two expressions to generate the correlation matrix $W$, as defined by Eq. 14, and then apply Eq. 15 and Eq. 9 to obtain $S_{vN}$ and $I_{A_{1}:A_{2}}$, respectively, for the long-time stationary state.
Numerical estimation of log negativity: Equation 11 suggests that to evaluate log negativity, the crucial part is to perform the partial transposition. Unfortunately, the partial transposition of a fermionic Gaussian state is not a Gaussian state [84; 85]. This makes it difficult to efficiently calculate log-negativity in non-interacting fermionic systems. However, by expressing the partial transposition as a linear combination of two Gaussian operators, an upper bound on log-negativity can be obtained and is expressed as [85; 84; 86; 87; 88; 40; 89]
@@
where trace norm of $O_{+}$ is given by
@@
where $\Gamma^{A}$ is the correlation matrix corresponding to the subsystem $A$ ($A=A_{1}\cup A_{2}$) and is defined as $\Gamma^{A}=2W-I$. With Eqs. 19, 21 and 22, the upper bound of log negativity can be evaluated using Eq. 18.
Figure 1: In the first row, we plot the mutual information between two subsystems, $I_{A_{1}:A_{2}}$, of the stationary state as a function of the post-quench chemical potential $\mu_{f}=\mu$, for three values of $\alpha$: $\alpha=0$ (black squares), $\alpha=1$ (magenta stars), and $\alpha=2$ (orange pluses). In the left column, the pre-quench chemical potential $\mu_{i}$ is set at the critical point, i.e., $\mu_{i}=1$, while in the right column it is set at a non-critical point, with $\mu_{i}=1.5$. In the second row, we plot the log-negativity of the stationary state using the same symbols and quench parameters as in the first row. In the third row, we plot the occupation probability $n_{k}$ of Bogoliubov fermions at the soft mode, calculated from Eq. 23, with the same symbols and quench parameters as in the first row. For all figures, the quench protocol for the strength of the $p$-wave pairing interaction $\Delta$ is chosen as $\Delta_{i}=-1$ and $\Delta_{f}=1$.

### v2

@@ Abstract
We investigate the signature of quantum criticality in the long-time stationary state of the long-range Kitaev chain by performing various quench protocols. In this model, the pairing interaction decays with distance according to a power law with exponent $\alpha$. Using quantum information-theoretic measures, such as mutual information and logarithmic negativity, we show that, irrespective of the values of $\alpha$, critical-to-critical quench displays quantum criticality even in the stationary state. Remarkably, in the presence of long-range pairing interactions, where fermionic correlators decay algebraically even at noncritical points, the signature of quantum criticality persists in the stationary state. Furthermore, the effective central charge, calculated from both mutual information and logarithmic negativity of the stationary state following a critical-to-critical quench, agrees with the central charge of the corresponding ground states for both $\alpha=0$ and $\alpha=2$. Therefore, information of the universality class can be inferred from the stationary state.
@@ I Introduction
Understanding quantum phase transitions (QPTs) has been a central topic of interest over the past few decades due to its significant implications for understanding the collective behavior of many-body systems [1]. QPTs have been both theoretically and experimentally studied extensively in various systems, including quantum spin chains [2; 3; 4; 5; 6; 7; 1; 8; 9; 10; 11; 12], Bose-Einstein condensates [13; 14; 15; 16; 17; 18; 19; 20], and strongly correlated electron systems [21; 22; 23; 24; 25]. QPT occurs at specific values of the parameter(s) present in the Hamiltonian, called critical points, where the system exhibits scale invariance. A universal behavior ensues due to the scale invariance, which can be classified into certain universality classes based on the space dimension and symmetry of the order parameter [26; 27; 28; 29]. This universality is studied using conformal field theories, and the corresponding central charge value is commonly used to identify the underlying universality class [30; 31; 32]. For instance, systems belonging to the Ising universality class are described by a minimal model with a central charge $c=1/2$, while the Luttinger liquid universality class corresponds to a conformal bosonic theory with $c=1$ [33; 34]. Note that at quantum critical points, the development of long-range correlations leads to an algebraic decay of correlators over distance, in contrast to the exponential decay typically observed at noncritical points in Hamiltonians with only short-range interactions [35; 36; 37; 38]. This is often used to identify critical points.
@@
While the existence of quantum criticality and long-range correlations in the ground states of many-body quantum systems are widely studied, their presence in stationary states is not yet well explored. In Ref. [39], it is shown that when two sides of a one-dimensional (1D) noninteracting fermionic chain with nearest-neighbor hopping are prepared at different temperatures, the mutual information scales logarithmically with subsystem size in the steady state, indicating the presence of long-range correlations. Another example of the survival of quantum criticality in the long-time stationary state is shown in the anisotropic $XY$ chain under a sudden quench protocol [40]. Using quantum information measures such as mutual information and logarithmic (log-) negativity, it is shown that when both the pre- and postquench parameters of the Hamiltonian are set at the critical point, both mutual information and log-negativity exhibit a peak, indicating the presence of criticality in the stationary state. In contrast, the peak in mutual information and log-negativity vanishes for other quench protocols when either/both pre- and postquench parameters differ from critical values. The signature of quantum criticality in the stationary state is attributed to a change in correlation pattern from exponential to algebraic decay, which was captured by mutual information and log-negativity.
@@
A natural question follows: What about the scenario when correlators decay algebraically, even at noncritical points? Long-range interacting systems provide an ideal platform to explore this question as fermionic correlators have been shown to decay algebraically over distance, even when the Hamiltonian is gapped [41; 42; 43; 44], i.e., at noncritical points. These long-range interacting systems have been studied in different contexts both theoretically and experimentally in atomic, molecular, and optical lattice systems [45; 46; 47; 48; 49; 50; 51; 52; 53; 54; 55; 56; 57; 58; 59; 60; 61]. We take the 1D long-range Kitaev model with a pairing term that decays with distance as a power law $\propto l^{-\alpha}$, with $\alpha$ being the exponent [44]. From an experimental perspective, this model is particularly relevant as it is closely related to the Ising model with tunable long-range interactions, which can be experimentally realized using trapped ion setups [53; 54; 55; 62; 58; 57]. In the ground state, this model undergoes an exotic transition from the Ising-type universality class, observed for $\alpha>3/2$, to a Luttinger liquid universality class at $\alpha=0$.
@@
In this paper, we investigate whether it is possible to capture quantum criticality in the stationary state when fermionic correlators decay algebraically even at noncritical points. Additionally, we examine whether the stationary state can be described by the same universality class as the ground state by analyzing the central charge obtained via scaling of mutual information and log-negativity for various quench protocols. Specifically, we consider three values of $\alpha$ corresponding to different universality classes observed in the ground state: first, corresponding to the Ising universality class ($\alpha=2$); second, the Luttinger liquid universality class ($\alpha=0$); and a third value ($\alpha=1$) where there is no universality [44]. To understand the tripartite information in the postquench stationary state and whether signatures of criticality will manifest, we study tripartite mutual information.
@@ II The model and ground state phase transition
where $f_{j}(f_{j}^{\dagger})$ is the fermionic annihilation (creation) operator at site $j$, satisfying the canonical anticommutation relations $\{f_{i},f_{j}^{\dagger}\}=\delta_{ij}$, $\{f_{i},f_{j}\}=\{f_{i}^{\dagger},f_{j}^{\dagger}\}=0$. The parameters $t$ and $\mu$ represent the tunneling rate between two neighboring sites and the chemical potential, respectively. We set $12t=$ 1 in the rest of the paper. The parameter, $\Delta$ denotes the strength of the fermion $p$-wave pairing interaction, while its range is governed by the exponent $\alpha\in[0,\infty)$. The two limits, i.e., $\alpha=0$ and $\alpha\to\infty$, correspond to all-to-all interaction with equal strength and nearest-neighbor interaction, respectively. We consider the antiperiodic boundary condition throughout the paper. The Hamiltonian in Eq. 1 is exactly solvable. To see this, we first perform a Fourier transformation to transform Eq. 1 in the momentum space as
@@
where $g_{\alpha}(k)=\sum_{l=1}^{N-1}\frac{\sin(kl)}{l^{\alpha}}$ and $k_{n}=\frac{2\pi}{N}\left(n+1/2\right)$. In the large-$N$ limit, $g_{\alpha}(k)$ takes the form
@@
The function $\rm Li_{\alpha}(z)$ represents the polylogarithm of the complex variable $z$ of order $\alpha$. The Hamiltonian in Eq. 2 can further be cast to a diagonal form by performing the following Bogoliubov transformation:
@@
with $\theta$ being the Bogoliubov angle, and defined as
@@
Equation (6) signifies that each mode is independent, implying the integrability of the system. In the limit $\alpha\to\infty$, only nearest-neighbor terms contribute to the sum in $g_{\alpha}(k)$, resulting in $g_{\infty}(k)=\sin(k)$. The Hamiltonian in Eq. 1 in this limit coincides with the Hamiltonian of the $XY$ model obtained via the Jordan-Wigner transformation. For the $XY$ model, a ground-state phase transition occurs from the gapped ferromagnetic ordered phase to the gapped paramagnetic disordered phase with increasing the chemical potential $|\mu|$, with the two critical gapless points at $\mu=\pm 1$ [44]. The spectrum in Eq. 7 continues to be gapless at $\mu=\pm 1$, even with the smaller values of $\alpha$, i.e., in the presence of long-range interactions as long as $\alpha>1$. However, the phase diagram changes for $\alpha<1$. This regime is referred to as “strong” long-range regime since $\alpha<d$, with $d$ representing the dimension of the system. In comparison to $\alpha>1$, one significant change is that $\mu=-1$ is no longer a critical point, and the phase diagram is no longer symmetric across the line $\mu=0$ [44; 42].
Table 1: Summary of the effective central charge $c_{\rm eff}$ for different values of $\alpha$ and $\mu$ for the ground state of the long-range Kitaev model [44; 42; 63; 64].
@@
A good probe to identify these critical points is the von Neumann entropy $S_{vN}$, expressed as $S_{vN}=-{\text{Tr}}\left(\rho_{L}\ln\rho_{L}\right)$, where $\rho_{L}$ is the reduced density matrix of the subsystem of length $L$. At these critical points, where the spectral gap closes, $S_{vN}$ for short-range interacting system scales as [65; 66; 67; 68]
@@
where $c$ is the central charge of the underlying conformal field theory (CFT), $N$ is the total system size, and $c^{\prime}$ is a nonuniversal constant. In fact, Eq. 8 is also valid for all values of $\alpha$ and $\mu$, but the central charge $c$ is now replaced with an effective central charge $c_{\rm eff}$ [44; 63]. The values of the $c_{\rm eff}$ for the ground state are summarized in Table 1. It is worth mentioning that $\alpha=0$ signifies all-to-all coupling between the sites and falls under the Tomonaga-Luttinger liquid universality class [69].
@@ III Methods
In the last section, we summarized criticality seen in the ground state of the LRK model and (effective) central charge of the underlying CFT. A natural question then arises: Does this criticality survive when the system undergoes out-of-equilibrium dynamics? In other words, in a stationary state, can we find any signature of quantum criticality? A common way to achieve the nonequilibrium dynamics is through sudden quench where the local or global parameters of the system are suddenly changed [70]. As a consequence, the ground state of the prequench Hamiltonian is no longer a ground state of the postquench Hamiltonian, but a superposition of its eigenstates. The long-time stationary state of the LRK model under sudden quench is the system of interest. To probe the criticality, much like the ground state, entanglement entropy would be our first choice. However, as the long-time stationary state typically involves highly excited states, $S_{vN}$ exhibits volume-law scaling eclipsing the logarithmic dependence on subsystem size which may be present due to criticality. So, to extract these logarithmic correlations, we consider quantum information-theoretic measures such as mutual information and log-negativity. For the ground state, mutual information has already been established to capture the criticality in fermionic systems [71].
Mutual information. The mutual information between two subsystems $A_{1}$ and $A_{2}$ is defined as
@@
where $S_{vN}^{A_{1}}(S_{vN}^{A_{2}})$ is the von Neumann entropy of the subsystem $A_{1}(A_{2})$ and $S_{vN}^{A_{1}\cup A_{2}}$ is the von Neumann entropy of $A_{1}\cup A_{2}$. Throughout the paper, we consider the subsystem sizes $|A_{1}|=|A_{2}|=L$. The mutual information between two subsystems measures the total amount of information (both classical and quantum) that one subsystem contains about the other [72]. Furthermore, if $S_{vN}$ of the long-time stationary state contains both the volume and logarithmic terms, the mutual information should scale as $I_{A_{1}:A_{2}}\sim b\ln L$. Thus, if the long-range correlations are present in the stationary state, then the mutual information should behave as [73; 74; 75; 76; 39; 40]
@@
Log-negativity.: Similar to mutual information, log-negativity, an entanglement monotone, also quantifies the nonlocal correlation between two subsystems. But unlike mutual information, it captures only quantum correlations [77; 78; 79; 80]. It is defined as [78; 77]
@@
where $c$ is the central charge. For the $XY$ model, during a critical-to-critical quench, the log-negativity of the long-time stationary state scales as $\xi_{A_{1}:A_{2}}=\frac{1}{8}\ln L$, where $L_{1}=L_{2}=L$ [40]. The logarithmic divergence of $\xi_{A_{1}:A_{2}}$ is consistent with the critical ground state, as the $XY$ model belongs to the Ising universality class with $c=1/2$. Based on this, we assume that if there is any signature of long-range quantum correlations in the long time stationary state, the log-negativity would exhibit the following scaling behavior:
@@
where $c_{\rm eff}^{N}$ represents the effective central charge of the stationary state extracted using the log-negativity.
Time evolution. For noninteracting free-fermionic chains, owing to the simple relationship between the eigenvalues of the reduced density matrix and two-point correlation function, we can calculate the entanglement entropy more efficiently by defining correlation matrix $W$ as
@@
$S=-\frac{1}{2}{\text{Tr}}\left[(I-W)\ln(I-W)+W\ln W\right].$ (15)
At any finite time $t$, the correlation functions are given by $C_{nm}(t)=\langle f_{n}^{\dagger}(t)f_{m}(t)\rangle$ and $F_{nm}(t)=\langle f_{n}(t)f_{m}(t)\rangle$, where the expectation values are taken with respect to the time evolved state obtained under the operation of the postquench Hamiltonian. The two time-dependent correlation functions, $C_{nm}(t)$ and $F_{nm}(t)$, are expressed as [83]
$\langle f_{j}(t)f_{j+l}(t)\rangle$ $=\frac{1}{2N}\sum_{n=0}^{N-1}e^{ik_{n}l}\bigg{(}-\sin\left[2\lambda^{f}_{\alpha}(k_{n})t\right]\sin\left(2\delta\theta_{k_{n}}\right)$ (16)
$\quad+i\bigg{\{}\sin\left(2\theta_{k_{n}}\right)\cos\left(2\delta\theta_{k_{n}}\right)$
$\qquad-\cos\left(2\lambda^{f}_{\alpha}(k_{n})t\right)\sin\left(2\delta\theta_{k_{n}}\right)\cos\left(2\theta_{k_{n}}\right)\bigg{\}}\bigg{)}$
$\langle f_{j}^{\dagger}(t)f_{j+l}(t)\rangle$ $=\frac{1}{2N}\sum_{n=0}^{N-1}e^{-ik_{n}l}\bigg{\{}1-\cos\left(2\theta_{k_{n}}\right)\cos\left(2\delta\theta_{k_{n}}\right)$
$\quad-\sin\left(2\theta_{k_{n}}\right)\sin\left(2\delta\theta_{k_{n}}\right)\cos\left(2\lambda^{f}_{\alpha}(k_{n})t\right)\bigg{\}},$
where $\delta\theta_{k_{n}}=\theta^{f}_{k_{n}}-\theta^{i}_{k_{n}}$ is the difference between pre- and postquench Bogoliubov angles defined in Eq. 5 and $\lambda^{f}_{\alpha}(k_{n})$ is defined by Eq. 7 for the postquench Hamiltonian.
In the limit $t\to\infty$, the time-dependent sine and cosine functions become highly oscillatory and the respective sum would tend to zero. Using this approximation, the expression for the correlation functions in the long-time stationary state is given by
@@
where $\langle f_{j}^{\dagger}f_{j+l}\rangle_{\text{st}}$ and $\langle f_{j}f_{j+l}\rangle_{\text{st}}$ are the two-point correlation and two-point anomalous correlation functions for the stationary state. We utilize the above two expressions to generate the correlation matrix $W$, as defined by Eq. 14, and then apply Eq. 15 and Eq. 9 to obtain $S_{vN}$ and $I_{A_{1}:A_{2}}$, respectively, for the long-time stationary state.
Numerical estimation of log-negativity. Equation (11) suggests that to evaluate log-negativity, the crucial part is to perform the partial transposition. Unfortunately, the partial transposition of a fermionic Gaussian state is not a Gaussian state [84; 85]. This makes it difficult to efficiently calculate log-negativity in noninteracting fermionic systems. However, by expressing the partial transposition as a linear combination of two Gaussian operators, an upper bound on log-negativity can be obtained and is expressed as [85; 84; 86; 87; 88; 40; 89]
@@
where the trace norm of $O_{+}$ is given by
@@
where $\Gamma^{A}$ is the correlation matrix corresponding to the subsystem $A$ ($A=A_{1}\cup A_{2}$) and is defined as $\Gamma^{A}=2W-I$. With Eqs. 19, 21 and 22, the upper bound of log-negativity can be evaluated using Eq. 18.
Figure 1: Top row: The mutual information between two subsystems, $I_{A_{1}:A_{2}}$, of the stationary state as a function of the postquench chemical potential $\mu_{f}=\mu$, for three values of $\alpha$: $\alpha=0$ (black squares), $\alpha=1$ (magenta stars), and $\alpha=2$ (orange pluses). The prequench chemical potential $\mu_{i}$ is set at the critical point (left column), i.e., $\mu_{i}=1$, and at a noncritical point (right column), i.e., $\mu_{i}=1.5$. For noncritical initial state (right column), blue triangles correspond to $\alpha=0$, red circles to $\alpha=1$ and green crosses to $\alpha=2$. Middle row: The log-negativity of the stationary state using the same symbols and quench parameters as the top row. Bottom row: The occupation probability $n_{k}$ of Bogoliubov fermions at the soft mode, calculated from Eq. (23), with the same symbols and quench parameters as in the top row. For all figures, the quench protocol for the strength of the $p$-wave pairing interaction $\Delta$ is chosen as $\Delta_{i}=-1$ and $\Delta_{f}=1$.